# Multiple Invertible and Partial-Equivariant Function for Latent Vector Transformation to Enhance Disentanglement in VAEs

### v1

@@ C.1 Integration for Multiple IPE-Transformation and EF-Conversion
$\begin{split}\mathcal{L}_{rec}&\coloneqq\frac{1}{k}\sum_{i=1}^{k}\Bigg{[}\int q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\mathsf{d}\hat{{\bm{z}}}_{i}\prod_{j=1,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\Bigg{]}\\ &=\frac{1}{k}\sum_{i=1}^{k}E_{q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}})),\end{split}$ (27)
@@
$\log p_{\theta}({\bm{x}})$ $=\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}})\mathsf{d}\hat{{\bm{z}}}^{\prime}$ (28)
$=\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log\frac{p_{\theta}({\bm{x}},\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\mathsf{d}\hat{{\bm{z}}}^{\prime}$ (29)
$\begin{split}&=\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\cdot\\ &~{}~{}~{}~{}\Big{[}\log\frac{p_{\theta}({\bm{x}},\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}-\log\frac{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\Big{]}\mathsf{d}\hat{{\bm{z}}}^{\prime}\end{split}$ (30)
$\geqq\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log\frac{p_{\theta}({\bm{x}},\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\mathsf{d}\hat{{\bm{z}}}^{\prime}$ (31)
$\begin{split}&=\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\cdot\\ &~{}~{}~{}~{}\Big{[}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})+\log\frac{p(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\Big{]}\mathsf{d}\hat{{\bm{z}}}^{\prime},\end{split}$ (32)
@@
$\begin{split}p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})&=\frac{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k}|{\bm{x}})p_{\theta}({\bm{x}})}{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})}\\ &=\frac{p_{\theta}({\bm{x}})\prod_{i=1}^{k}p_{\theta}(\hat{{\bm{z}}}_{i}|{\bm{x}})}{\prod_{i=1}^{k}p_{\theta}(\hat{{\bm{z}}}_{i})}(\because(\hat{{\bm{z}}}_{i}\mathrel{\perp\!\!\!\perp}\hat{{\bm{z}}}_{j}s|{\bm{x}}))\\ &=\prod_{i=1}^{k}\frac{p_{\theta}(\hat{{\bm{z}}}_{i}|{\bm{x}})p_{\theta}({\bm{x}}^{\frac{1}{k}})}{p_{\theta}(\hat{{\bm{z}}}_{i})}\\ &=p_{\theta}({\bm{x}})^{-(k-1)}\prod_{i=1}^{k}\frac{p_{\theta}(\hat{{\bm{z}}}_{i}|{\bm{x}})p_{\theta}({\bm{x}})}{p_{\theta}(\hat{{\bm{z}}}_{i})}\\ &=p_{\theta}({\bm{x}})^{-(k-1)}\prod_{i=1}^{k}p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i}),\end{split}$ (33)
where $\hat{{\bm{z}}}_{j}s=\cap_{j=1,j\neq i}^{k}\hat{{\bm{z}}}_{j}$. Therefore,
@@
$\small\begin{split}&\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})\mathsf{d}\hat{{\bm{z}}}^{\prime}\\ &=\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\Big{[}-(k-1)\log p_{\theta}({\bm{x}})+\prod_{i=1}^{k}p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\Big{]}\mathsf{d}\hat{{\bm{z}}}^{\prime}\\ &=-(k-1)\log p_{\theta}({\bm{x}})+\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\prod_{i=j}^{k}p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{j})\mathsf{d}\hat{{\bm{z}}}^{\prime}.\end{split}$ (34)
Then,
@@
$\begin{split}\log p_{\theta}({\bm{x}})&\geq\frac{1}{k}\sum_{i=1}^{k}\Bigg{[}\int q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\mathsf{d}\hat{{\bm{z}}}_{i}\prod_{j=1,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\Bigg{]}\\ &-\int\prod_{i}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log\frac{\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})}{\prod_{i=1}^{k}p(\hat{{\bm{z}}}_{i})}\mathsf{d}\hat{{\bm{z}}}^{\prime}\\ &=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q(\hat{{\bm{z}}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\\ &-\sum_{i=1}^{k}\bigg{[}D_{\mathrm{KL}}(q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})||p(\hat{{\bm{z}}}_{i}))\prod_{j=1,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\bigg{]}\\ &=\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})-\sum_{i=1}^{k}D_{\mathrm{KL}}(q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})||p({\bm{z}}_{i}))\\ &=\frac{1}{k}\Big{[}\sum_{i=1}^{k}\mathbb{E}_{q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})-kD_{\mathrm{KL}}(q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})||p(\hat{{\bm{z}}}_{i}))\Big{]}.\end{split}$ (35)
@@
However, following Eq. 36, k samples are generated, and each sample is disentangled for different factors. We implement the output as the average of the sum of the k samples to obtain a single sample with a superposition effect from k samples. Moreover, the KL divergence term in Eq. 36 represents that increasing number of MIPE-transformation is equal to increasing $\beta$ hyper-parameter in $\beta$-VAE Higgins et al. (2017).
@@ Appendix D Details of Experimental Environment
D.1 Model Architecture
@@
Table 8: VAE architecture for 3D Shapes, and 3D Cars datasets. For exceptional case, CLG-VAE, we ues ten dimension size on 3D Shapes dataset Zhu et al. (2021).
@@ D.2 Details of Setting
D.3 Impact of Symmetric Matrix Exponential
@@
D.4 Additional Experiment of Computing Complexity
@@
D.5 Qualitative Analysis
[…]
In Figure 8, $\beta$-TCVAE struggles with y-position and rotation, as shown on the 6th row, and with scale and shape represented on the 7th row. On the contrary, MIPET-$\beta$-TCVAE separates y-position and rotation factor (10th, and 7th rows), also the activated dimensions of MIPET-$\beta$-TCVAE are not overlapped with each factor. Applied our method on $\beta$-TCVAE shows better disentangled representation on dSprites dataset. These results also show that our proposed method improves disentangled representation learning. As shown in the Figure 8, $\beta$-VAE struggles with rotation and scale factors in 4thdimension. Also, it struggles with x-position and scale factors in 8th dimension, and x-position and rotation factors in 9th dimension. However, MIPET-$\beta$-VAE only struggles with rotation and shape factors in 5th dimension. As shown in the Figure 8, CLG-VAE struggles with rotation and shape factors in 2nd dimension, and shape and scale factors in 7th dimension. However, MIPET-CLG-VAE separates rotation and shape factors in 10th, and 1st dimensions respectively.
The qualitative analysis with 3D Shapes dataset, as shown in the Figure 11, $\beta$-TCVAE struggles with shape and scale in $5^{th}$ dimension. However MIPET-$\beta$-TCVAE As shown in the Figure 11, CLG-VAE struggles with shape and wall color factors in 4th dimension, and shape and object color factors in 7th dimension. In particular, it struggles with tree factors in 9th dimension. On the other hand, MIPET-CLG-VAE separates shape, wall, and object color factors.
The qualitative analysis with 3D Cars dataset, as show in Figure 11, the left side is the $\beta$-TCVAE result, and it struggles with body, and azimuth factors shown in the 7th row. However, MIPET-$\beta$-TCVAE separates azimuth (6th row) and body (1st row). In particular, MIPET-$\beta$-TCVAE learns color factor (3rd row) which does not exist on $\beta$-TCVAE.
Figure 6: Qualitative results on dSprites. The left-side grids are input images and their variants by changing activations of each dimension of latent vectors. The first row shows input images. The right-side table shows matching pre-defined factors of the dataset (red: MIPET, blue: no MIPET).
Figure 7: Qualitative analysis result of $\beta$-VAE and MIPET-$\beta$-VAE.
Figure 8: Qualitative analysis result of CLG-VAE (0.2) and MIPET-CLG-VAE (0.2) with dSprites.
Figure 9: The Shape is object shape, Orien is an orientation of object, Scale is a scale factor of object, Wall is wall color factor, Floor is floor color, and Object is object color factors. It represents the $\beta$-VAE ($\beta=2$) results.
Figure 10: Qualitative analysis result of CLG-VAE (0.2) and MIPET-CLG-VAE (0.2) with 3D Shapes.
Figure 11: Qualitative analysis result of $\beta$-VAE (4.0) with 3D Cars.

### v3

@@ Appendix C Objective Function
$\begin{aligned} \mathcal{L}_{rec}&\coloneqq\frac{1}{k}\sum_{i=1}^{k}\Bigg[\int q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\mathsf{d}\hat{{\bm{z}}}_{i}\prod_{j=1,\,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\Bigg]\\ &=\frac{1}{k}\sum_{i=1}^{k}E_{q_{\phi,\psi_{i}}({\bm{z}}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\psi_{i}({\bm{z}})),\end{aligned}$
@@
$\log p_{\theta}({\bm{x}})=\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}})\mathsf{d}\hat{{\bm{z}}}^{\prime},$ (29)
@@
$\begin{aligned} \log p_{\theta}({\bm{x}})\geqq\int\prod_{i=1}^{k}q_{1}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log\frac{p_{\theta}({\bm{x}},\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\mathsf{d}\hat{{\bm{z}}}^{\prime}.\end{aligned}$
Expanding Eq. 30 then
@@
$\begin{aligned} \log p_{\theta}({\bm{x}})&\geqq\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\Big[\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})\\ &~~~+\log\frac{p(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})}{q(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k}|{\bm{x}})}\Big]\mathsf{d}\hat{{\bm{z}}}^{\prime}.\end{aligned}$
The first term of Eq. 31 is expanded as:
$p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})=\frac{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k}|{\bm{x}})p_{\theta}({\bm{x}})}{p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})}.$ (32)
Each IPE-transformation function operates independently, then all $\hat{{\bm{z}}}_{i}$ are independent and conditionally independent given $\bm{x}$}. Then we have
$p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})=\frac{p_{\theta}({\bm{x}})\prod_{i=1}^{k}p_{\theta}(\hat{{\bm{z}}}_{i}|{\bm{x}})}{\prod_{i=1}^{k}p_{\theta}(\hat{{\bm{z}}}_{i})},$ (33)
where $\prod_{i=1}^{k}p_{\theta}(\hat{{\bm{z}}}_{i}|{\bm{x}})=p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k}|{\bm{x}})$ and $\prod_{i=1}^{k}p_{\theta}(\hat{{\bm{z}}}_{i})=p_{\theta}(\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})$. Let replace $p_{\theta}({\bm{x}})$ as $p_{\theta}({\bm{x}})^{-(k-1)}\prod_{i=1}^{k}p_{\theta}({\bm{x}})$. Then we have
$p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\ldots,\hat{{\bm{z}}}_{k})=p_{\theta}({\bm{x}})^{-(k-1)}\prod_{i=1}^{k}p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i}),$ (34)
Plugging Eq. 34 into the first term of Eq. 31, then we have
$\begin{aligned} &\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{1},\hat{{\bm{z}}}_{2},\cdots,\hat{{\bm{z}}}_{k})\mathsf{d}\hat{{\bm{z}}}^{\prime}\\ &=-(k-1)\log p_{\theta}({\bm{x}})+\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\Big[\sum_{j=1}^{k}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{j})\Big]\mathsf{d}\hat{{\bm{z}}}^{\prime}.\end{aligned}$
Applying same process between the Eq. 32 and Eq. 33, and plugging Eq. 35 into Eq. 31, then we have
$\begin{aligned} k\log p_{\theta}({\bm{x}})&\geqq\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\Big[\sum_{j=1}^{k}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{j})\Big]\mathsf{d}\hat{{\bm{z}}}^{\prime}\\ &~~~+\int\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log\frac{\prod_{i=1}^{k}p(\hat{{\bm{z}}}_{i})}{\prod_{i=1}^{k}q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})}\mathsf{d}\hat{{\bm{z}}}^{\prime}.\end{aligned}$
Expanding the first term and inverse the second term of Eq. 36, then we have
$\begin{aligned} \log p_{\theta}({\bm{x}})&\geq\frac{1}{k}\sum_{i=1}^{k}\Bigg[\int q_{i}(\hat{{\bm{z}}}_{i}|{\bm{x}})\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})\mathsf{d}\hat{{\bm{z}}}_{i}\prod_{j=1,\,j\neq i}^{k}\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}\Bigg]\\ &\quad-\sum_{i=1}^{k}D_{\mathrm{KL}}\left(q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})\,\|\,p({\bm{z}}_{i})\right)\end{aligned}$
Replacing $\int q_{j}(\hat{{\bm{z}}}_{j}|{\bm{x}})\text{d}\hat{{\bm{z}}}_{j}$ as $1$, then we have:
$\begin{aligned} \log p_{\theta}({\bm{x}})\geq\frac{1}{k}\Bigg[\sum_{i=1}^{k}\mathbb{E}_{q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})}\log p_{\theta}({\bm{x}}|\hat{{\bm{z}}}_{i})-k\,D_{\mathrm{KL}}\left(q_{\phi}(\hat{{\bm{z}}}_{i}|{\bm{x}})\,\|\,p(\hat{{\bm{z}}}_{i})\right)\Bigg]\end{aligned}$
@@
However, following Eq. 39, k samples are generated, and each sample is disentangled for different factors. We implement the output as the average of the sum of the k samples to obtain a single sample with a superposition effect from k samples. Moreover, the KL divergence term in Eq. 39 represents that increasing number of MIPE-transformation is equal to increasing $\beta$ hyper-parameter in $\beta$-VAE (Higgins et al., 2017).
@@ Appendix D Details of Experimental Environment
D.1 Model Architecture and Hyper-Parameter Tuning
The details of model architecture is demonstrated in Table 7and 7. The details of hyper-parameter settings are in Table 5(f).
@@ D.1 Model Architecture and Hyper-Parameter Tuning
Table 7: VAE architecture for 3D Shapes, and 3D Cars datasets. For exceptional case, CLG-VAE, we use ten dimension size on 3D Shapes dataset (Zhu et al., 2021).
@@
D.2 Details of Datasets
We benchmark MIPET-VAEs against several standard VAEs on the following datasets: 1) dSprites (Matthey et al., 2017) which consists of 737,280 binary $64\times 64$ images of dSprites with five independent ground truth factors(number of values), $i.e.$ shape(3), orientation(40), scale(6), x-position(32), and y-position(32). 2) 3D Shapes (Burgess and Kim, 2018) which consists of 480,000 RGB $64\times 64\times 3$ images of 3D Shapes with six independent ground truth factors: shape(4) orientation(15), scale(8), wall color(10), floor color(10), and object color(10). 3) 3D Cars (Reed et al., 2015) which consists of 17,568 RGB $64\times 64\times 3$ images of 3D Shapes with three independent ground truth factors: car models(183), azimuth directions(24), and elevations(4).
D.3 Details of Training and Evaluation
Training
We set 256 mini-batch size in the datasets (dSprites, 3D Shapes, and 3D Cars), Adam optimizer with learning rate $4\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and epochs from $\{30,67,200\}$ as a common setting for all the comparative methods. For the comparison, we follow training and inference on the whole dataset. We train each model for 30, 67, and 200 epochs on the dSprites, 3D Shapes, and 3D Cars, respectively, as introduced in Kim and Mnih (2018); Ren et al. (2022). We tune $\beta$ from $\{1,2,4,10\}$ and $\{4,6\}$ for $\beta$-VAE and $\beta$-TCVAE, respectively. We apply the proposed method to $\beta$-TCVAE only with the EF similarity loss term because $\beta$-TCVAE penalizes the divided KL divergence terms. We set the same encoder and decoder architecture in each model to exclude the overlapped effects. The $\alpha=1$ and $\gamma=1$ of $\beta$-TCVAE as Chen et al. (2018). We set the dimension size of the latent vectors from $\{6,10\}$ for 10 on dSprites and 3D Cars datasets and 6 for 3D Shapes, but we set 10 for CLG-VAE because it sets 10 dimensions size on 3D Shapes in Zhu et al. (2021). Regarding the CLG-VAE, we fix $\lambda_{\text{decomp}}$, $\lambda_{\text{hessian}}$, and forward group features as 40, 20, and 0.2, respectively. Because the hyper-parameters showed the best result in Zhu et al. (2021). We set group reconstruction from $\{0.2,0.5,0.7\}$. For Control-VAE, we set the maximum KL divergence value from $\{10,12,\ldots,20\}$. In addition, we set masking ratio $\lambda$ from $\{0.0,0.5,\ldots 2.0,\infty\}$. To check the impact of MIPE-transformation, we do not consider the Groupified VAE because the latter is implemented with an extended decoder (different capacity).
Evaluation
We conduct experiments on NVIDIA A100, RTX 2080 Ti, and RTX 3090. We set 100 samples to evaluate global empirical variance in each dimension and run it a total of 800 times to estimate the FVM score introduced in Kim and Mnih (2018). For the MIG (Chen et al., 2018), SAP (Kumar et al., 2018), and DCI (Eastwood and Williams, 2018), we follow default values introduced in Michlo (2021), training and evaluation 100 and 50 times with 100 mini-batches, respectively. We evaluate four disentanglement metrics for a less biased understanding of the actual states of disentanglement.
dataset Metrics
FVM MIG SAP DCI
D.4 Additional Result of Relation Between Factors and Latent Vector Dimensions
Figure 6: 3D Car Dataset: Azim. refers to the azimuth factor, and Elev. refers to the elevation factor.
In the 3D Car dataset, a trade-off was observed. While the maximum value across models and evaluation factors showed a slight decrease, the azimuth factor exhibited a notable improvement compared to the declines seen in other factors.
D.5 Impact of Symmetric Matrix Exponential
@@
D.6 Additional Experiment of Computing Complexity
@@
D.7 Semantic Mask: Mapping Gaussian Distribution to Diverse Exponential Family
As shown in Fig. 7, MIPET-VAE with semantic mask preserves the dataset structure compared to VAE. Also, EF conversion with semantic mask maps the Gaussian distribution to the exponential family compared to MIPET without the semantic mask and VAE cases.
Figure 7: 2D Exponential Family Distribution Datasets